Find financial statements for Intel (INTC) and Duke Energy (DUK). Which firm has the higher ratio of market value to book value of equity? Does this make sense to you? Why or why not? Which firms pays out a higher percentage of its profits as dividends to shareholders? Does this make sense - why or why not?
Answer #1

Market value to book value of equity = (Price per share * no. of shares outstanding)/(Total assets - total liabilities as reported in the balance sheet)

= (Price per share * no. of shares outstanding)/(book value of equity)

For both companies the latest annual report is available for the year ended 2018 and hence my computations are based on that.

Intel’s market value to book value of equity = ($211*4,611 million shares)/$74,563 million

= 13.05

Duke Energy’s market value to book value of equity = ($151 * 708 million shares)/$43,834 million

= 2.44

Thus Intel has the higher ratio of market value to book value of equity and this makes sense as Intel will have a higher fair value/market value given the fact that its operations is spread across the world and its products are used in many computing hardware across the world. On the other hand Duke Energy’s operation is an electric power holding company with operations mainly in America.

% of profits as dividends = dividend per share/earnings per share

For Intel the % of profits that it pays as dividends = 1.20/4.57 = 26.26%

For Duke Energy the % of profits that it pays as dividends = 3.64/3.76 = 96.81%

Thus Duke Energy has a much higher dividend payout ratio and this makes sense because utilities companies generally pay a large portion of their earnings in the form of dividend to shareholders.